### v3

# LLM-Based Detection of Tangled Code Changes for Higher-Quality Method-Level Bug Datasets

Abstract.
Tangled code changes, commits that conflate unrelated modifications such as bug fixes, refactorings, and enhancements, introduce significant noise into bug datasets and adversely affect the performance of bug prediction models. Addressing this issue at a fine-grained, method-level granularity remains unexplored. This is critical to address, as recent bug prediction models, driven by practitioner demand, are increasingly focusing on finer granularity rather than traditional class- or file-level predictions. This study investigates the utility of Large Language Models (LLMs) for detecting tangled code changes by leveraging both commit messages and method-level code diffs. We formulate the problem as a binary classification task and evaluate multiple prompting strategies, including zero-shot, few-shot, and chain-of-thought prompting, using state-of-the-art proprietary LLMs such as GPT-5 and Gemini-2.0-Flash, and open-source models such as GPT-OSS-120B and CodeBERT.
Our results demonstrate that combining commit messages with code diffs significantly enhances model performance, with the combined few-shot and chain-of-thought prompting achieving an F1-score of 0.883. Additionally, we explore machine learning models trained on LLM-generated embeddings, where a multi-layer perceptron classifier achieves superior performance (F1-score: 0.906, MCC: 0.807). Applying our approach to 49 open-source projects improves the distributional separability of code metrics between buggy and non-buggy methods, demonstrating the promise of LLMs for method-level commit untangling and potentially contributing to improving the accuracy of future bug prediction models.
Tangled code changes, large language models, machine learning, code diffs, commit messages, prompting techniques
copyright: acmlicensedjournalyear: 2025conference: X; Month from–to, 2025; X, Y
1. Introduction
Software maintenance often exceeds the cost of initial development (Börstler and Paech, 2016), with bug detection and resolution alone consuming 50–70% of total development expenses (wang_bugpre_2023). Undetected bugs can result in cascading failures, reduced developer productivity, and increased technical debt (christine fisher, 2020; noauthor_most_2024). Consequently, the research community has extensively investigated bug prediction techniques to facilitate timely and targeted maintenance interventions (Chowdhury et al., 2024; pascarella_performance_2020; wattanakriengkrai_predicting_2022; Giger et al., 2012; rahman_relationships_2017). Early bug prediction models focused on class- or file-level granularity (zimmermann_predicting_2007; Alsolai et al., 2018; Basili et al., 1996; Bell et al., 2011; Gil and Lalouche, 2017; sun_boosting_2025; yin_line-level_2025). However, these models are less helpful in practice, as only a small fraction of code in large units is typically defective (Mo et al., 2022; pascarella_performance_2020). Line-level models offer finer granularity but are prone to false positives due to coincidental line similarity (Chowdhury et al., 2024; servant_fuzzy_2017; Grund et al., 2021). This has driven growing interest in method-level bug prediction (Chowdhury et al., 2024; Mo et al., 2022; pascarella_performance_2020; Shippey; Giger et al., 2012; Hata et al., 2012; Menzies et al., 2007), which aligns more closely with real-world debugging practices (Grund et al., 2021; Chowdhury et al., 2024). However, method-level bug prediction faces challenges including evaluation bias and noisy data (pascarella_performance_2020; Chowdhury et al., 2024), and progress is constrained by the scarcity of clean, labeled datasets (Chowdhury et al., 2024; pascarella_performance_2020).
A key obstacle in building clean bug datasets is the presence of noise in training data (Bird et al., 2009), often caused by tangled code changes (Chowdhury et al., 2024; PârΩ, 2020; Herzig et al., 2016). While Version Control Systems like Git are intended for atomic commits, developers frequently combine unrelated tasks, such as bug fixes, refactorings, and enhancements, into a single commit (Herzig and Zeller, 2013; Dias et al., 2015; Kirinuki et al., 2014), violating best practices and introducing tangled changes. Researchers typically use commit messages to identify bug-fix commits, assuming all associated modifications are relevant. However, this assumption fails in the presence of tangled commits, leading to false positives in labels, which degrades dataset quality and impairs bug prediction performance (Herzig et al., 2016; Herbold et al., 2022).
To mitigate the impact of tangled changes, several prior works have proposed various untangling techniques ranging from heuristic-based program slicing (Muylaert and De Roover, 2018), lexical and data-flow based models (PârΩ, 2020), and graph-based learning using dependency and name flows (Li et al., 2022; xu_detecting_2025). While effective at the file or statement level, none of these methods have addressed the problem at the method-level granularity, leaving a gap that presents a unique opportunity for exploration. Method-level untangling cannot be achieved using file-level approaches, as a single file may contain multiple methods, only some of which are bug-related. Labeling all methods in a file as buggy is inappropriate and would produce numerous false positives. On the other hand, although statements can be aggregated to form method-level representations, determining whether an individual statement contributes to a bug fix requires understanding its broader context, including surrounding code, method semantics, and commit intent (tufano2019empirical; sintaha2023katana). For developers and models alike, the relevant context may span from within the method to the entire file or even across files (sintaha2023katana). For instance, in the refill function of FixedIntervalRateLimiter.java from Apache HBase Commit 84a5039  [^1] https://github.com/apache/hbase/commit/84a50393ee56d09abb68f54b44b64f5279bd33c9, the function signature modification appears to be a bug fix when viewed in isolation, as removing an argument typically alters the function’s logic; however, examining the full method reveals it is actually a refactoring change. Thus, simply applying a statement-level model repeatedly at the method-level would require redundant context analysis for each modified line, introducing unnecessary computational overhead. Therefore, distinct representations at different granularity are needed to effectively model bug-fix behavior while maintaining contextual relevance.
Given that, this study focuses on detecting tangled changes at the method-level by framing the task as a binary classification problem: determining whether a method-level code change is related to a bug fix (Buggy) or not (NotBuggy) based on commit messages and code diffs. To tackle this, we explore the use of Large Language Models (LLMs), which have demonstrated strong capabilities in analyzing natural language and source code (Chew et al., 2023; Nam et al., 2024). In contrast to prior approaches in code analysis, classification, and untangling that rely on language-specific algorithms (Herzig et al., 2016; Guo and Song, 2017) or complex models requiring extensive training and feature engineering (zeng2024colare), LLM-based methods enable faster system development through their language-agnostic design. LLMs also provide advantages such as promptability, flexibility in input handling, and the ability to leverage contextual signals from both code and commit metadata.
We evaluate several prompting strategies using both proprietary and open-source LLMs. Additionally, we develop machine learning models using LLM generated embeddings. Finally, we evaluate the impact of the LLM-based technique by measuring improvements in the distributional separability of code metrics between Buggy and NotBuggy methods through statistical analysis. In general, the contribution of the paper is centered around four research questions.
RQ1: Can zero-shot LLMs detect tangled code changes by using code diff and commit message?
We designed a classification-based approach that uses LLMs to detect tangled changes from method-level code diffs, using zero-shot prompting, with or without the corresponding commit messages. Through experiments on our curated gold dataset, we show that including the commit message significantly improves the performance across multiple LLM variants, achieving the highest F1-score of 0.879 using gpt-5.
RQ2: How do different prompting techniques influence the effectiveness of detecting tangled code changes?
We investigate the impact of advanced prompting strategies, such as few-shot, chain-of-thought, and a hybrid combination. Our analysis shows that LLMs can reason more accurately about change semantics when guided with structured prompts and examples. Among these, chain-of-thought + few-shot prompting with gpt-4o achieves the most balanced performance (F1-score: 0.883) and demonstrates superior capability in handling complex inputs.
RQ3: How well can embedding-based machine learning models detect tangled code changes?
We utilized LLMs to generate embeddings from commit messages and code diffs, which served as input for embedding-based classifiers. One of the models exhibits strong predictive capabilities, outperforming the results obtained in RQ1 and RQ2. Specifically, a Multi-layer Perceptron classifier achieved the highest F1-score of 0.906, demonstrating the effectiveness of embedding-based representations for detecting tangled changes.
RQ4: What is the potential impact of LLM-based untangling on future method-level bug prediction models?
Using a Less-Noisy dataset, created by filtering out noisy samples through our LLM-based approach, we observe that the distributional differences in various code metrics between Buggy and NotBuggy methods are significantly more than in the original noisy dataset. This finding shows promise for future machine learning (ML) models for bug prediction, as many of them rely on these code metrics to predict whether a method is bug-prone or not.
To enable replication, we share our data and code publicly. [^2] https://github.com/SQMLab/Tangled
2. Related Works
This section reviews prior work on tangled code changes, their prevalence, impact, and untangling techniques, as well as LLM-based approaches for analyzing and classifying text and code. Together, these studies motivate our research.
2.1. Tangled Changes
Herzig and Zeller (Herzig and Zeller, 2013) first used the term Tangled Changes and reported that up to 15% of bug-fixing commits across several open-source Java projects contained tangled changes. Subsequently, Tao and Kim (tao_partitioning_2015) later observed tangling in up to 29% of revisions. Herbold et al. (Herbold et al., 2022) confirmed its prevalence and impact, showing that only 22–38% of lines in bug-fix commits actually fix bugs. Kochhar et al. (Kochhar et al., 2014) further found that 28% of files in bug-fix commits contain no bugs.
The prevalence of tangled changes greatly affects the historical analysis of source code. By labeling all the modified methods in a bug-fixing commit as buggy methods, it harms the bug prediction models significantly (Herzig et al., 2016; Chowdhury et al., 2024). Consequently, researchers focused on detecting and untangling commits (Herzig and Zeller, 2013; tao_partitioning_2015; Guo and Song, 2017; Muylaert and De Roover, 2018; PârΩ, 2020).
Early techniques relied on static analysis and heuristics. Herzig et al. (Herzig and Zeller, 2013) introduced confidence voters based on file distance, change coupling, call graphs, and data dependencies to estimate task co-membership. Tao and Kim (tao_partitioning_2015) applied program slicing and pattern matching to cluster semantically related edits, achieving 69% agreement with manual decompositions. Guo and Song (Guo and Song, 2017) proposed CHGCUTTER, an interactive method using control and data dependencies to decompose composite changes while preserving syntactic correctness. Muylaert and De Roover (Muylaert and De Roover, 2018) applied program slicing on abstract syntax trees to untangle composite commits by grouping fine-grained code changes according to their dependence graph slices. Pârtachi et al. (PârΩ, 2020) introduced Flexeme, which overlays name flows on program dependency graphs using a $\delta$-NFG structure to better capture lexical cues. Using Agglomerative clustering for commit untangling they developed a tool Heddle, which achieved improved clustering accuracy (F1-score 0.81) and runtime efficiency.
Graph-based modeling has emerged as a powerful approach for capturing complex code relationships. Shen et al. (shen_smartcommit_2021) proposed SmartCommit, representing diff hunks as graph nodes enriched with semantic links (hard, soft, refactoring, cosmetic), and achieved a decomposition accuracy between 71-84%. Chen et al. (Chen et al., 2022) and Xu et al. (xu_detecting_2025) enhanced this approach by incorporating node attributes such as token content and control/data flow, achieving 7.8% and 8.2% higher accuracy than Flexeme, respectively.
Fan et al. (Fan et al., 2024) introduced a Heterogeneous Directed Graph Neural Network designed to capture semantic dependencies without relying on explicit code links. By leveraging hierarchical graphs at the entity and statement levels, they achieved substantial untangling improvements, 25% for C# and 19.2% for Java, outperforming Flexeme and SmartCommit without sacrificing time efficiency.
Recent work increasingly applies machine learning to learn untangling patterns from data. UTANGO (Li et al., 2022) uses a Graph Convolutional Network to learn contextual embeddings of code changes, incorporating cloned code and surrounding context. Framed as supervised clustering, it outperforms Flexeme by 9.9% accuracy. Dias et al. (Dias et al., 2015) and Liu et al. (Liu et al., 2018) explored fine-grained tracking and clustering based on developer interactions and commit metadata to achieve high temporal and contextual resolution.
Although these methods have shown success at the file- or statement-level, the method-level remains unexplored, despite being preferred by practitioners and researchers in order to create a clean and noise-free method-level bug dataset (pascarella_performance_2020; Chowdhury et al., 2024). Motivated by this gap, we investigate whether tangled changes can be detected from method-level diffs and their corresponding commit messages.
2.2. LLMs for Code Analysis & Classification
LLMs have shown strong performance in natural language understanding and source code analysis (Brown et al., 2020; zheng2024few). Trained on large corpora of text and code, they generalize well to unseen inputs and often perform effectively in zero-shot settings (Kojima et al., 2022). This versatility has fueled growing interest in using LLMs for text classification and code analysis (Ahmad et al., 2021; Feng et al., 2020). Building on these capabilities, our study explores LLMs for detecting tangled code changes by jointly analyzing commit messages and code diffs.
Prompt design is critical to optimizing LLM performance. Recent advances in prompt engineering have introduced zero-shot, few-shot, and chain-of-thought prompting (wei_finetuned_2022; Brown et al., 2020; wei_chain--thought_2023). Zero-shot learning uses only task instructions, while few-shot prompting includes exemplar input-output pairs to guide responses, yielding improved performance across tasks (Brown et al., 2020). Chain-of-thought prompting enhances reasoning by encouraging step-wise problem decomposition (wei_chain--thought_2023). Motivated by their empirical success, we systematically evaluate these techniques for detecting tangled code changes.
LLMs are built on Transformer architectures, consisting of encoder and decoder modules (vaswani_attention_2017). Encoders convert input text into high-dimensional embeddings that capture semantic relationships, making them effective for downstream classification (petukhova_text_2024; Keraghel et al., 2024). We extend this utility to software engineering by generating embeddings from commit messages and code diffs, and training machine learning models to classify them as tangled or untangled.
LLM embeddings are also used in commit classification tasks (zeng2025first; zeng2024colare), which classifies commits into maintenance activity types. CodeBERT (Feng et al., 2020) has been used in COLARE (zeng2024colare) to classify commits into corrective, adaptive, and perfective categories by integrating hunk-level code representations with commit messages and file features. However, such models are build for commit-level classification which does not properly align with our method-level objectives.
In summary, this study is motivated by three factors: (1) the lack of fine-grained research on the tangled changes at the method level, especially for bug prediction (Herzig et al., 2016; Chowdhury et al., 2024); (2) the proven capabilities of LLMs in analyzing textual and code artifacts via advanced prompting (Brown et al., 2020; zheng2024few; wei_finetuned_2022; wei_chain--thought_2023); and (3) the semantic richness of LLM-generated embeddings for classification (petukhova_text_2024; Keraghel et al., 2024). To the best of our knowledge, this is the first study to evaluate LLMs for detecting tangled code changes within bug-fix commits at the method level.
3. Methodology
Figure 1 shows the overview of our approach that we use to evaluate LLMs in distinguishing Buggy diff and NotBuggy diff. In this section, the components of the methodology are described step by step.
Figure 1. The methodology for RQ1 and RQ2 consists of three parts: (a) dataset, (b) prompt preparation, and (c) LLM Query. The prompt setup for few-shot + chain-of-thought prompting consists of four instruction parts: (i) persona, (ii) task description, (iii) behavioral guidance, and (iv) output formatting. For Diff-only detection of RQ1, references to the commit message are excluded from everywhere. Other experiments use the full instructions, with the output formatting modified to require a single-word answer in non-chain-of-thought setups. Few-shot examples are included only in few-shot experiments.:
3.1. Dataset
To assess whether LLMs can be used to untangle real bug fix changes from other changes, a dataset is required that includes both genuine bug fix code diffs and non-bug-fix code diffs. We provide the model with a code diff with or without the commit message to evaluate its ability to distinguish between bug-fix and non-bug-fix changes.
Fortunately, an existing dataset by Chowdhury et al. (Chowdhury et al., 2024) contains 774,051 Java methods from 49 open source projects, where each method is associated with its complete change history and is enriched with metadata such as commit messages, timestamps, authorship, and change type. The dataset also includes code diffs for all historical changes, labels indicating whether each change is related to a bug fix or not, and information on the number of methods modified within the same commit.
However, the dataset contains noise. In some cases, methods involved in a bug fix commit may not contribute to the actual fix, representing instances of tangled changes. We use this dataset to construct a curated gold dataset by removing such noise. The construction process involves a two-step sampling strategy, combining automated filtering and manual labeling, to ensure high quality and accurate representation of tangled change scenarios.
Automated Approach. The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) includes the total number of methods modified in each commit. They recommended that if a method is the only one modified in a bug fix commit, it can be confidently considered a bug-related change. We classify such code diffs as Buggy without ambiguity. Using this criterion, we collected 730 Buggy diffs. They also suggested that a method can be considered NotBuggy if it has never appeared in a bug fix commit throughout its lifetime. Based on this guideline and leveraging the large pool of eligible NotBuggy methods, we randomly selected 730 samples that have no history of being part of any bug fix commit. During this process, three duplicate diffs, likely caused by duplicated code across codebases, were identified and removed, resulting in 727 unique NotBuggy instances. Thus, the Automated step of our gold dataset construction produced a total of 1,457 method-level examples. We intentionally limited the number of NotBuggy samples, as querying proprietary LLMs incurs significant computational and financial costs. Including all such samples would have significantly increased the overall cost.
Manual Approach. While the Automated approach ensured high confidence in labeling, it did not capture the complexity of real-world tangled change scenarios, where multiple unrelated modifications often occur within a single commit. To address this limitation, we introduced a second step focused on incorporating tangled examples. In this step, we extracted additional method-level diffs from bug-fix commits that modified multiple methods. The first author, a graduate student with over two years of industry experience in software engineering, manually labeled each method-level change to determine whether it was related to a bug fix (Buggy) or was incidental, such as refactorings or unrelated enhancements (NotBuggy) Each method was reviewed in conjunction with its associated commit message to evaluate semantic alignment with bug-fixing intent. The evaluation emphasized structural corrections, error handling logic, and domain-specific repairs. To promote annotation consistency and minimize errors, we developed a user interface that allowed side-by-side inspection of code diffs and their corresponding commit metadata. Using this manual approach, we identified 166 tangled Buggy methods and 141 tangled NotBuggy methods from 450 randomly selected commits in the dataset of Chowdhury et al. (Chowdhury et al., 2024).
Combining all steps, the gold dataset consists of 1,764 method-level change instances, each labeled as either Buggy or NotBuggy, based on their actual relevance to bug fixing activities. To evaluate labeling reliability, an independent validation was conducted by another graduate student in software engineering, who randomly picked 100 samples from the manually labeled portion of the dataset and labeled them independently. Some disagreements arose, mainly due to differing interpretations of what constitutes a bug. Specifically, the first labeler marked some quality-related fixes as NotBuggy, while the second considered them Buggy. Following a discussion with the last author, quality-fix-related changes were finalized as NotBuggy. The inter-rater agreement, measured using Cohen’s Kappa (Cohen, 1960), was 0.82, indicating strong consistency and confirming the reliability of the labeling process. Consequently, this dataset provides a robust benchmark for both training and evaluating our detection approaches, while also offering a valuable foundation for future research in this domain.
3.2. Prompt Preparation
The design of our prompt structure draws on recent advancements in prompt engineering that demonstrate the effectiveness of decomposing complex instructions into clearly scoped components (Brown et al., 2020; wei_chain--thought_2023; zhou_least--most_2023). As prior work has shown, LLMs are highly sensitive to prompt formulation, and performance can vary substantially based on how the task, instructions, and examples are framed (Liu et al., 2023). To maximize reliability and adaptability, we designed a modular prompt template consisting of three core components, aligned with prior studies (Amatriain, 2024; Giray, 2023), as illustrated in Figure 1.
Figure 2. Commit message and two methods modified in the same commit: the message indicates a bug fix, but only the method on the right contains the bug-fix change, while the method on the left is a non-bug-related tangled change. This was used as few-shot examples in our prompts.:
Instructions. This component defines the LLM’s role and behavior. It is composed of four parts: (i) a role-based persona (e.g., “You are a Git commit review assistant”), which helps constrain the model’s responses by grounding its behavior and improves performance (Hu and Collier, 2024); (ii) a high-level task description (Amatriain, 2024); (iii) step-by-step behavioral guidance to structure the model’s reasoning (wei_chain--thought_2023); and (iv) explicit output formatting requirements to reduce ambiguity and improve parsability (Liu et al., 2024).
Few-shot Examples. Few-shot prompting has been shown to significantly improve performance by enabling in-context learning, especially in code and software engineering tasks (zheng2024few). For few-shot examples, as shown in Figure 2, we selected a representative bug-fix commit containing one diff genuinely related to a bug fix and another unrelated change, a real instance of tangled code changes. For the negative example (NotBuggy), we selected the code diff that merely replaced constant values with equivalent boolean literals, without altering the program’s logic or behavior. In contrast, the other code diff (Buggy) addressed an issue in the byte-to-hexadecimal conversion logic by correctly handling signed bytes using bit masking, rectifying a functional error. These two fixed examples were consistently utilized to prime the LLMs in all experiments employing few-shot learning. These examples were selected to represent contrasting scenarios, one clearly indicative of a bug-fix and the other representing a refactor or performance-related change that does not qualify as a bug-fix, although both of the examples come from the same commit.
LLMs often exhibit significant performance degradation as prompt length increases, due to limitations in handling extended context (Levy et al., 2024; zhang_inftybench_2024). Additionally, longer prompts incur higher computational and monetary costs, making token efficiency a practical concern. Since method-level diffs can vary greatly in length and the input query, over which we have no control, is the primary driver of prompt size, we deliberately restricted the prompt to a single example per class (Buggy and NotBuggy). This prompting approach helps maintain prompt lengths within manageable token limits across different model families. By limiting the number of in-context examples, we strike a balance between performance and resource usage. This strategy aligns with established practices in few-shot prompting  (Brown et al., 2020), where concise and diverse instances are used to anchor the model’s reasoning prior to inference.
Query. This includes the method-level code diff and its corresponding commit message.
Each research question required slight variations in prompt configuration, e.g., enabling or disabling few-shot examples, toggling chain-of-thought reasoning, or isolating input modalities (diff-only vs. diff + message). All corresponding prompt templates are included in our replication package, located within the prompts directory.
3.3. Selecting and Querying LLMs
Selecting LLMs. For RQ1 and RQ2, we selected six LLMs: four proprietary, GPT-4o-mini, GPT-4o, GPT-5, and Gemini-2.0-Flash and two open-source, GPT-OSS-20B and GPT-OSS-120B. The proprietary LLMs were chosen based on their performance on recent code-related benchmarks (Google DeepMind, 2024; OpenAI, 2025; Batista et al., 2024; Bruni et al., 2025), robust API support, and cost-effectiveness for large-scale experimentation. GPT-4o serves as OpenAI’s full-scale model offering superior robustness over its smaller variant, GPT-4o-mini (OpenAI, 2025; Bruni et al., 2025), while GPT-5 represents the latest generation of OpenAI’s LLMs. Gemini-2.0-Flash, from Google (Google DeepMind, 2024), delivers competitive performance with the GPT series and has demonstrated effectiveness in code-related tasks (Batista et al., 2024).
For open-source alternatives and to ensure reproducibility and accessibility, we choose GPT-OSS-20B (requiring  16 GB GPU memory) and GPT-OSS-120B (runnable on a single 80 GB GPU) (Agarwal et al., 2025). Despite their moderate scale, these models deliver reasoning and code-understanding performance comparable to proprietary baselines (Agarwal et al., 2025; Bi et al., 2025), balancing efficiency and capability.
For RQ3, we employed four models for embedding generation: one proprietary and three open-source. The proprietary model, Gemini’s text-embedding-004 offers high-quality embeddings optimized for classification tasks (Lee et al., 2024). Open-source models, CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020) are widely adopted for code representation learning (Ma et al., 2024), while EmbeddingGemma-300m, a recent lightweight model from Google DeepMind, is designed for low-latency, high-throughput applications (vera2025embeddinggemma). The models used in RQ1 and RQ2 were not included in RQ3 because they are decoder-only generative models optimized for text generation, whereas RQ3 required encoder-based architectures specialized for producing fixed-length embeddings suitable for classification tasks.
In summary, proprietary models were selected for their state-of-the-art performance across benchmarks, including code-related tasks (Batista et al., 2024; Bae et al., 2024), while open-source models were selected for their competitive accuracy and efficiency relative to larger open-source alternatives.
Querying LLMs. The three parts of the prompt (instruction + examples + query) are concatenated and formatted according to the input constraints of the respective LLM. The final query prompt is submitted to the LLM, which classifies the change as either Buggy or NotBuggy. Outputs are reviewed and filtered automatically, if required, to ensure alignment with evaluation criteria.
3.4. Evaluation Metrics
We evaluate model performance using standard classification metrics, including accuracy, precision, recall, and F1-score. Additionally, we use the Matthews Correlation Coefficient (MCC), a robust metric that considers all four outcomes: true positives, true negatives, false positives, and false negatives. Unlike accuracy, MCC provides a more balanced measure. An MCC of 1 indicates perfect prediction, 0 indicates random guessing, and –1 reflects total disagreement between predictions and actual labels.
3.5. Experimental Setup
All experiments involving open-source LLMs were conducted on a High-Performance Computing (HPC) cluster equipped with Intel 8570 CPUs (2.1 GHz) and multiple NVIDIA H100 GPUs. The number of CPUs, GPUs and the memory allocation varied by model, depending on resource availability and scheduling within the cluster.
4. Approach, Analysis & Results
This section answers the four research questions and outlines the specific methodologies used to investigate each one.
4.1. RQ1: Performance of zero-shot LLMs using code diff and commit message
Both the commit message and the corresponding code diff contain information about the nature of the code changes. While the diff captures the structural modifications to the source code, the commit message provides a natural language description of the developer’s intent. LLMs offer a powerful mechanism to detect semantic signals indicative of bug-fixing behavior across these modalities. By leveraging their ability to reason over structured and unstructured data (code + natural language) (Nam et al., 2024; Chew et al., 2023), LLMs offer a promising alternative to traditional rule-based or static analysis techniques. Furthermore, using prompt-based inference avoids the need for task-specific fine-tuning, making the approach adaptable across repositories and languages. However, the effectiveness of LLMs is sensitive to input length, as inference cost increases with the number of tokens in the prompt. Given that both code diffs and commit messages can vary significantly in length, it becomes essential to evaluate whether both are necessary for effective detection or whether the code diff alone suffices. This research question seeks to empirically determine whether commit messages provide additive value beyond the code diff for detecting tangled changes.
We investigate two scenarios using LLMs by designing tailored prompting strategies for the binary classification of method-level diffs. Since no examples are provided within the prompt, this setup qualifies as zero-shot prompting.
Diff-Only. The LLM is provided only with the code diff and instructed to assess whether the change represents a bug fix (Buggy) or not (NotBuggy). The prompt is adapted from Figure 1, with all references to commit messages and few-shot examples removed, and the instruction modified to require a single-word response.
Diff+Message. The prompt includes both the code diff and its commit message, instructing the model to assess their semantic alignment to identify bug-fixing intent. This prompt follows the Diff-Only Detection setting, except that all references to commit messages are retained as in Figure 1.
Table 1 compares the performance of the evaluated LLMs under two input settings: Diff-only and Diff+Message. In the Diff-only scenario, gpt-5 achieves the best overall performance, yielding the highest F1-score (0.767) and MCC (0.490). When commit messages are included, all models show substantial improvements. The gpt-5 again attains the highest F1-score (0.879), while gpt-4o provides the best accuracy (0.874), precision (0.868) and MCC (0.747). Although gpt-5 achieves the top F1-score, its high precision–recall gap indicates less balanced performance, whereas gpt-4o offers the most consistent results across all metrics.
The gemini-2.0-flash and the gpt-oss models achieve high recall (>0.9>0.9) but low precision (<0.76<0.76) resulting in reduced F1-scores. The gpt-oss models perform remarkably well in Diff+Message scenario despite their smaller size narrowing the gap with proprietary models (within  3–4% of gpt-5’s F1-score), likely benefiting from their unique dual-channel chat format (Agarwal et al., 2025), where an internal “analysis” channel generates intermediate reasoning (chain-of-thought) before producing the answer on ”final” channel. Overall, these results demonstrate that integrating commit messages with diffs significantly improves performance, and that open-source LLMs are also effective despite their smaller sizes.
Table 1. Performance comparison of LLMs for Diff-only and Diff+Message. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”). For each scenario, the best value for each metric is highlighted in bold.
Scenario
LLM
Accuracy
Precision
Recall
F1-score
MCC
Diff-only gpt-5 0.742 0.709 0.834 0.767 0.490
gpt-4o-mini 0.704 0.660 0.860 0.747 0.426
gpt-4o 0.691 0.654 0.831 0.732 0.395
gemini-2.0-flash 0.692 0.650 0.853 0.738 0.402
gpt-oss-20b 0.663 0.615 0.900 0.730 0.364
gpt-oss-120b 0.667 0.619 0.900 0.733 0.372
Diff
Message
gpt-5 0.871 0.839 0.924 0.879 0.746
gpt-4o-mini 0.822 0.764 0.941 0.843 0.661
gpt-4o 0.874 0.868 0.886 0.877 0.747
gemini-2.0-flash 0.796 0.717 0.988 0.831 0.639
gpt-oss-20b 0.820 0.758 0.950 0.843 0.662
gpt-oss-120b 0.827 0.766 0.948 0.847 0.671
Summary of RQ1: Both proprietary and open-source LLMs can effectively detect tangled code changes using code diffs, with or without commit messages, though combining both inputs yields better results than code diffs alone.
4.2. RQ2: Performance of different prompting techniques
The effectiveness of LLMs highly depends on the input prompt, and the success of different prompting techniques is proven in various studies (wei_finetuned_2022; Brown et al., 2020; wei_chain--thought_2023). This research question investigates how different prompting techniques, namely few-shot learning, chain-of-thought prompting, and a hybrid approach combining both, impact the classification performance of LLMs.
Few-shot. The prompt includes labeled examples of buggy and non-buggy changes to guide the model, while the instruction remains identical to the Diff+Message setting of RQ1.
Chain-of-thought. The model is instructed to explain its reasoning in the output. The exact instruction prompt is shown in Figure 1, and no examples are included.
Chain-of-thought + Few-shot. This setting combines the above two techniques by including labeled examples and requiring the model to explain its reasoning in the output.
Based on the findings from RQ1, we use the combination of commit message and code diff for RQ2. Additionally, we exclude gpt-4o-mini from further experiments, as gpt-4o significantly outperformed this small model. This decision also supports cost efficiency, as the prompting techniques used in RQ2 involve substantial token consumption in both input and output.
Table 2 presents the performance of different prompting strategies across LLMs. While gpt-4o with few-shot achieves the highest precision (0.903) but lower recall (0.833), other models exhibit the opposite trend. In this setting, gpt-5 attains the best overall performance with an F1-score of 0.884. For chain-of-thought prompting, performance generally decreases across models, though gpt-5 again achieves the highest F1-score (0.872). When combining few-shot and chain-of-thought prompting, gpt-4o delivers the best results across all metrics except recall, which is highest for gemini-2.0-flash.
Table 2. Performance comparison of different prompting techniques. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Prompting Technique
LLM
Accuracy
Precision
Recall
F1-score
MCC
Few-shot gpt-5 0.876 0.842 0.931 0.884 0.757
gpt-4o 0.870 0.903 0.833 0.866 0.742
gemini-2.0-flash 0.802 0.723 0.989 0.835 0.649
gpt-oss-20b 0.837 0.775 0.958 0.857 0.694
gpt-oss-120b 0.819 0.753 0.959 0.843 0.663
Chain
of
thought
gpt-5 0.863 0.829 0.920 0.872 0.730
gpt-4o 0.827 0.770 0.940 0.846 0.669
gemini-2.0-flash 0.765 0.688 0.984 0.810 0.587
gpt-oss-20b 0.829 0.770 0.946 0.849 0.676
gpt-oss-120b 0.829 0.765 0.956 0.850 0.679
Chain
of
thought
Few-shot
gpt-5 0.863 0.829 0.921 0.873 0.731
gpt-4o 0.880 0.871 0.896 0.883 0.760
gemini-2.0-flash 0.807 0.731 0.979 0.837 0.651
gpt-oss-20b 0.844 0.795 0.932 0.858 0.697
gpt-oss-120b 0.816 0.752 0.952 0.840 0.655
Among all prompting strategies, gpt-5 with few-shot yields the top F1-score, while gpt-4o attains the highest accuracy, precision, and MCC with chain-of-thought+few-shot. Considering balanced performance across metrics, particularly the small gap between precision and recall (0.025), gpt-4o with chain-of-thought+few-shot offers the most balanced results, with an F1-score (0.883) nearly identical to the best (0.884). Open-source models also benefit consistently, though modestly, from advanced prompting; notably, gpt-oss-20b remains competitive, often matching or surpassing gpt-oss-120b.
A comparison between Table 1 and Table 2 reveals that few-shot prompting improves performance primarily for gpt-5 and slightly for gpt-oss-20b, while other models show marginal or decreased performance. Notably, chain-of-thought alone results in poorer performance across most models. However, combining few-shot prompting with chain-of-thought yields the most balanced and consistent outcomes overall, as observed with gpt-4o. This trend aligns with prior research suggesting that few-shot prompting can lead models to overfit to example formats rather than the task itself, while zero-shot chain-of-thought prompting may amplify biases and degrade reasoning quality (Kojima et al., 2022; shaikh_second_2023). Our findings thus corroborate earlier studies indicating that integrating few-shot and chain-of-thought generally outperforms zero-shot approaches (Kojima et al., 2022).
Table 3. Predictions on a semantically misleading commit using chain-of-thought + few-shot prompting with gpt-4o.
Commit Message Diff (Excerpt) Ground Truth Prompting Method Prediction
Fixed bug 1050173 – ImmutableFieldRule no longer reports false positives for static fields. Also fixed version number in PMD.java.
- NameOccurrence occurance = …
- if (occurance.isOnLeftHandSide())
+ NameOccurrence occ = …
+ if (occ.isOnLeftHandSide())
NotBuggy Commit msg + diff (zero-shot) Buggy
Few-shot Buggy
Chain-of-thought NotBuggy
Chain-of-thought + Few-shot NotBuggy
4.2.1. Case analysis: an example of the effectiveness of chain-of-thought approach in a semantically ambiguous commit
To better understand why the chain of thought technique improves model performance in the few shot setting with gpt-4o, we analyze a representative example involving semantic ambiguity between the commit message and the code modifications from the PMD  [^3] https://github.com/pmd/pmd project. As shown in Table 3, the commit message clearly describes a bug fix involving two files, PMD.java and ImmutableFieldRule.java. In this commit [^4] https://github.com/pmd/pmd/commit/a405d23dfb9e574e2b2ef23f1f45d548a738ed3b, multiple methods are modified, including initializedInConstructor() from ImmutableFieldRule.java, which only includes variable renaming (e.g., occurance to occ) without affecting the program’s logic or behavior. No structural fix is present that addresses the reported issue. Despite this, both the zero-shot and few-shot models incorrectly classify the change as Buggy, influenced by the commit message. In contrast, the chain-of-thought technique, with or without few-shot prompting, correctly identifies the change as NotBuggy, consistent with the manually assigned label.
The reasoning produced by the chain-of-thought prompt illustrates a more grounded assessment. An important part of the reasoning states: ”. . . However, in the provided diff, the only change is renaming the variable occurance to occ, which is a simple refactor for readability or consistency purposes. This modification itself doesn’t address the logic or behavior of the code in terms of resolving the stated bug. The commit message also mentions a version number fix in PMD.java, but no relevant changes are shown in this diff. . . .”
This case highlights the effectiveness of chain-of-thought in systematically analyzing complex commit messages that reference multiple issues. While such messages offer valuable context, they can also introduce ambiguity. The step-by-step reasoning decomposed the message, independently assessed each part against the code diff, and correctly identified the lack of structural or semantic alignment, ultimately concluding that the change was not a bug fix.
Summary of RQ2: Combining chain-of-thought reasoning with few-shot prompting yields the most balanced and consistent performance across models. While few-shot prompting alone benefits mainly gpt-5, and chain-of-thought alone often reduces accuracy, their combination enhances reasoning stability and mitigates overfitting. This hybrid approach also improves interpretability, enabling models to correctly handle semantically misleading commits through more context-aware reasoning.
4.3. RQ3: Performance of embedding-based machine learning models
This research question explores the viability of embedding-based ML models for detecting tangled code changes. Unlike prompt-based approaches that rely on in-context reasoning of LLMs, this method uses embeddings, vector representations of the input data from commit messages and code diffs, to train supervised classifiers. The key objective is to assess whether these embeddings, when used as features in ML models, can achieve competitive performance in detecting bug-related code changes. This experiment is motivated by the previous studies where embedding-based models performed well and even outperformed prompt-based techniques (petukhova_text_2024; Keraghel et al., 2024).
The methodology employed for this research question diverges from that of the preceding research questions and the process illustrated in Figure 1. Specifically, this experiment omits the complex prompt engineering. Instead, only the core query, comprising the code diff and commit message, is retained and directly input to the LLM to generate fixed-length vector embeddings for downstream classification. The modified methodology is shown in Figure 3.
Figure 3. The methodology for RQ3 has two main parts: (i) embedding generations and (ii) building ML models.:
To generate embeddings, Gemini’s text-embedding-004 was employed with the classification task type. Due to resource constraints, this study focuses exclusively on this proprietary model, utilizing Gemini’s free quota. For the open-source models, a sliding-window approach with mean pooling was applied to address token input limits, 2048 for EmbeddingGemma-300m and 512 for CodeBERT and GraphCodeBERT. All models produced 768-dimensional vector representations for each instance, derived from the concatenation of the code diff and its corresponding commit message.
The classification task was treated as a binary problem, and the dataset was split into an 80% training set and a 20% test set. We evaluated multiple classifiers, including RandomForest (RF), Support Vector Classifier (SVC), DecisionTree (DT), and Multi-Layer Perceptron (MLP). These models were selected to represent a diverse range of learning paradigms: tree-based ensembles (RF, DT), margin-based optimization (SVC), and neural networks (MLP), which have been widely applied in various software engineering tasks (pascarella_performance_2020; Chowdhury et al., 2024).
Table 4. Performance comparison of ML models trained on text-embedding-004 embeddings. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Classifier Accuracy Precision Recall F1-score MCC
RF 0.816 0.841 0.82 0.83 0.629
SVC 0.884 0.918 0.866 0.891 0.768
DT 0.686 0.717 0.706 0.712 0.366
MLP 0.895 0.929 0.876 0.902 0.791
Table 4 presents the performance comparison of all models trained on text-embedding-004 embeddings, showing that the MLPClassifier achieved the best results, with an accuracy of 0.895, F1-score of 0.902, and MCC of 0.791, outperforming other classifiers in all metrics. Similar experiments were conducted using open-source embeddings; however, due to space limitations, only the MLP results are reported in Table 5, as it consistently outperformed other classifiers. Among the open-source models, GraphCodeBERT combined with MLP achieved the highest performance (F1-score = 0.831), exhibiting a modest gap of around 7% compared to the proprietary model. This suggests that while proprietary embeddings retain a advantage, open-source alternatives offer competitive and cost-effective performance for this task.
Table 5. Performance comparison of MLP classifier trained on different LLM embeddings. The first row of text-embedding-004 is taken from Table 4 for comparison. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Model Accuracy Precision Recall F1-score MCC
text-embedding-004 0.895 0.929 0.876 0.902 0.791
EmbeddingGemma 0.80 0.815 0.82 0.818 0.593
CodeBERT 0.765 0.738 0.833 0.782 0.533
GraphCodeBERT 0.828 0.831 0.83 0.831 0.656
Furthermore, we evaluated the MLPClassifier using the text-embedding-004 embeddings under a Leave-One-Out (LOO) strategy to test its robustness, where a single example is used as the test sample and the remaining serve as the training data. This approach not only offers more robust evaluation, but also can show if a model’s accuracy can be improved with more training samples. Encouragingly, the model achieved improved performance with an accuracy of 0.9036, precision of 0.8980, recall of 0.9141, F1-score of 0.906, and MCC of 0.8073, setting a new benchmark compared to prior best results in RQ1 and RQ2. In terms of class-wise performance the model maintained a balanced detection capability, with precision and recall exceeding 0.89 for both classes and achieving F1-scores of 0.91 for Buggy and 0.9 for NotBuggy.
Summary of RQ3: Embedding-based ML models, particularly MLPClassifier, demonstrate strong predictive power for tangled change detection, achieving over 90% F1-score and outperforming prior baselines of RQ1 and RQ2. More encouragingly, our results suggest that the performance of embedding-based models can be improved with an enlarged training dataset.
4.4. RQ4: Potential impact of LLM-based untangling on future method-level bug prediction models
From the results of RQ1 to RQ3, we observe that the LLM-based method-level tangled change detection approach performs well with high accuracy and F1-scores using both prompting techniques and embedding-based ML models. This suggests that such approaches can effectively reduce noise in bug datasets. However, an important question remains: Does this noise reduction benefit machine learning-based bug prediction models? Building such a model, however, requires addressing several challenges. First, a broad set of code metrics is needed, whereas the dataset from Chowdhury et al. (Chowdhury et al., 2024) includes only five. Second, Mashhadi et al. (Mashhadi et al., 2023) demonstrate that combining code metrics with embeddings leads to better performance, requiring us to generate embeddings for all methods. Finally, a thorough evaluation across multiple algorithms and fine-tuning strategies is essential. Due to these requirements, constructing a full bug prediction model is beyond the scope of this paper. Nonetheless, if we can demonstrate that our noise reduction approach improves the ability of code metrics to distinguish between bug-prone and non-bug-prone methods, likely, the performance of machine learning models relying on these metrics will also improve. Therefore, this research question examines whether the distribution of code metrics becomes more distinguishable in the dataset produced by our method.
In RQ1 and RQ2, we evaluated multiple prompting strategies and LLMs, showing that GPT-4o combined with few-shot + chain-of-thought prompting delivers the best and balanced performance comparing all metrics. In RQ3, we examined the feasibility of using LLM-generated embeddings with downstream machine learning classifiers to detect tangled changes. Although the latter approach yields better results, it requires generating embeddings for all tangled code diffs from bug-fix commits to support prediction with the machine learning model, which is more complex and time-consuming than the few-shot + chain-of-thought approach. Therefore, we employed GPT-4o with few-shot + chain-of-thought prompting to detect tangled changes across all methods in 49 software projects in a fast and efficient manner.
In the dataset by Chowdhury et al. (Chowdhury et al., 2024), each method is associated with its change history and the number of other methods modified in the same commit. We apply the few-shot and chain-of-thought strategy using GPT-4o to separate NotBuggy changes within bug-fix commits, thereby removing noise. To enable replication, the algorithm for creating the Less-Noisy dataset is included in the shared repository (algorithm/algorithm.png). After constructing the Less-Noisy dataset, we obtain four types of methods:
Noisy NotBuggy. Following Chowdhury et al. (Chowdhury et al., 2024), methods are labeled as NotBuggy if they are at least two years old and never involved in a bug-fix. This set is called Noisy because it comes from the original noisy dataset, which excludes non-buggy methods that were mixed with buggy changes.
Less-Noisy NotBuggy. It expands on the previous set by including methods confirmed as NotBuggy after LLM-based untangling.
Noisy Buggy. Following Chowdhury et al. (Chowdhury et al., 2024), methods are labeled as Buggy if they appear in bug-fix commits, without accounting for tangling, hence the dataset is considered Noisy.
Less-Noisy Buggy. This set includes methods identified as truly buggy after applying our LLM-based untangling approach.
The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) contains five code metrics computed for each version of a method across its history. For our analysis, we consider the metrics from the first version of each method to ensure consistency across the dataset. These code metrics are defined as follows: Size refers to the number of source lines of code, excluding comments and blank lines (Chowdhury et al., 2022, 2024); Readability is a score that reflects the ease of reading the source code, developed by Buse et al. (Buse and Weimer, 2009); McCabe measures cyclomatic complexity by counting the number of independent execution paths in a method (McCabe, 1976); FanOut represents the number of distinct methods called by a given method, indicating its dependency footprint (pascarella_performance_2020; Chowdhury et al., 2024); and Maintainability Index (MI) is a composite metric that combines several aspects, including complexity and size, to offer a comprehensive assessment of code maintainability (Oman and Hagemeister, 1992).
Figure 4. Distribution of code metrics for Buggy and NotBuggy methods in the Titan project across the Noisy and Less-Noisy datasets. Metric distributions differ more significantly in the Less-Noisy dataset.:
Figure 4 presents the distribution of these 5 code metrics across the four method sets for the project Titan. Visually, the difference in Size between Less-Noisy Buggy and Less-Noisy NotBuggy methods appears larger than the other two sets. A similar pattern is observed for all other code metrics. To determine whether these differences are statistically significant, we use the Wilcoxon rank-sum test (McKnight and Najab, 2010). This non-parametric test is widely adopted in software engineering research due to its robustness in comparing distributions (Bangash et al., 2020; Chen et al., 2020; Chowdhury et al., 2019). The results show that the differences in distributions are indeed statistically significant ($p\leq 0.05$) for all code metrics.
As the differences are statistically significant for both datasets, we calculated Cliff’s delta effect size (Macbeth et al., 2011) to understand the magnitude of these differences. This non-parametric measure is also widely used in software engineering research (Bangash et al., 2020; Chen et al., 2020; Chowdhury et al., 2019) and does not require assumptions about the data distribution. The results show that, in the Noisy dataset, the effect size is large for MI and medium for all other code metrics. In the Less-Noisy dataset, the effect size is large across all code metrics.
To evaluate the generalizability of our results, we performed Wilcoxon rank-sum tests across all 49 projects. In the Noisy dataset, significant differences ($p\leq 0.05$) were observed in 47 projects (95.92%) for Size, McCabe, and MI, 45 projects (91.84%) for Readability, and 48 projects (97.96%) for FanOut. In the Less-Noisy dataset, we observe improved results, with one additional project exhibiting statistically significant separation for each metric.
To further quantify the differences between Buggy and NotBuggy methods, we compute Cliff’s Delta effect sizes across all 49 projects. Table 6 presents the effect size distributions for both the Noisy and Less-Noisy datasets. We categorize the effect sizes following Hess et al. (Hess and Kromrey, 2004): Negligible (<0.147<0.147), Small ($0.147\leq\delta<0.33$), Medium ($0.33\leq\delta<0.474$), and Large ($\delta\geq 0.474$).
Table 6. Cliff’s Delta effect size values compare Noisy and Less-Noisy datasets, grouped by effect size category and shown as percentages in the format Noisy / Less-Noisy. The percentages for the Noisy dataset do not sum to 100% because the RxJava project was excluded as it had no NotBuggy methods.
Metric Negligible Small Medium Large
Size 0 / 0 14.29 / 6.12 46.94 / 20.41 36.73 / 73.47
Readability 2.04 / 2.04 36.73 / 24.49 42.86 / 40.82 16.33 / 32.65
McCabe 2.04 / 2.04 34.69 / 8.16 36.73 / 42.86 24.49 / 46.94
FanOut 0 / 0 10.20 / 6.12 46.94 / 22.45 40.82 / 71.43
MI 0 / 0 12.24 / 6.12 34.69 / 16.33 51.02 / 77.55
The table reports the proportion of projects falling into each effect size category for all five code metrics. The results indicate that, for most metrics, the Less-Noisy dataset shows a substantial shift toward larger effect sizes. For example, for the code metric, Size, in the Less-Noisy dataset, 73% of the projects exhibit a large effect size in the difference between Buggy and NotBuggy methods, compared to only 36% in the Noisy dataset. Similar patterns are observed for the other four code metrics. These findings clearly demonstrate that our filtering approach effectively eliminates noise and increases the distinction between the code metrics of buggy and non-buggy methods. As a result of this enhanced separability, we believe that our approach will support improved performance in future machine learning models for bug prediction.
Summary of RQ4: After constructing the Less-Noisy dataset using the LLM-based approach, statistical analysis reveals significantly improved separability (with larger effect sizes) in code metrics between buggy and non-buggy methods. This suggests that the noise reduction could enhance the performance of future ML-based bug prediction models.
5. Threats to Validity
Construct Validity. The manual annotation of code diffs as Buggy or NotBuggy may be affected by subjective bias. Although annotators were provided with explicit guidelines and a UI tool to assist in their evaluations, subjective judgment and semantic ambiguity in commit messages may have influenced labeling outcomes.
External Validity. The dataset used in this study comprises only open-source Java projects, which may limit the generalizability of our findings. Applying this approach to other programming languages, project domains, or commit conventions may require modifications to the prompts or retraining of models.
Internal Validity. The dataset used in this study was curated from method histories of Java projects obtained through CodeShovel (Grund et al., 2021), which is not fully accurate. Any inaccuracies in the method histories could impact the results.
Conclusion Validity. It is affected by all of the above mentioned threats.
6. Conclusion and Future Work
Method-level bug prediction is considered one of the holy grails in software engineering research. However, it remains an open challenge, primarily due to the absence of a noise-free method-level bug dataset (Chowdhury et al., 2024; pascarella_performance_2020). To address this, we evaluated the effectiveness of various LLMs and prompting strategies for detecting tangled changes at the method level. Our results show that even zero-shot LLMs achieve high accuracy when both commit messages and code diffs are provided (RQ1). Performance improves further when combining chain-of-thought with few-shot prompting (RQ2). Additionally, embedding-based models deliver even greater accuracy gains (RQ3). Across RQ1 to RQ3, open-source models also demonstrate strong performance. While they do not yet match proprietary models, they offer efficient alternatives with acceptable trade-offs. Building on these insights, we generated a Less-Noisy dataset using our LLM-based method. This dataset shows promise in developing more accurate method-level bug prediction models as it exhibits a significantly stronger power of commonly used code metrics to differentiate between bug-prone and non-bug-prone methods compared to the original Noisy dataset (RQ4).
The implications of our findings extend beyond the untangling problem and contribute to broader discussions on the capabilities and limitations of LLMs in software engineering. Although challenges such as hallucinations, reproducibility, and data leakage remain (yao_llm_2024; Ji et al., 2023; ye_cognitive_2023; sallou_breaking_2024), our study offers empirical evidence supporting the effectiveness of LLMs in tasks such as code reasoning, classification, and dataset construction (zheng2024few; Jaoua et al., 2025; Nam et al., 2024). In particular, our work shows that when guided by carefully designed prompts, LLMs can significantly improve software engineering workflows.
Looking ahead, a promising direction for future research is to enrich the dataset with additional code metrics and embeddings and to explore a range of machine learning models for building more effective method-level bug prediction models. To support continued progress, we have publicly released our code and dataset, allowing researchers to use them directly or adapt them to their specific needs.